# Capsule Vision 2024 Challenge: Multi-Class Abnormality Classification for Video Capsule Endoscopy

### v2

@@ Abstract
We present the Capsule Vision 2024 Challenge: Multi-Class Abnormality Classification for Video Capsule Endoscopy. It was virtually organized by the Research Center for Medical Image Analysis and Artificial Intelligence (MIAAI), Department of Medicine, Danube Private University, Krems, Austria, and Medical Imaging and Signal Analysis Hub (MISAHUB) in collaboration with the 9th International Conference on Computer Vision & Image Processing (CVIP 2024) being organized by the Indian Institute of Information Technology, Design and Manufacturing (IIITDM) Kancheepuram, Chennai, India. This document provides an overview of the challenge, including the registration process, rules, submission format, description of the datasets used, qualified team rankings, all team descriptions, and the benchmarking results reported by the organizers.
@@ VI-A Description of all the received entries
Team PuppyOps utilized DINOv2, combining a vision transformer for feature extraction and fully connected layers for multi-class classification. Random flips and brightness adjustments were applied as data augmentation techniques.
@@
Team MedInfoLab IIT Hyderabad performed fine-tuning of BiomedCLIP-PubMedBERT by integrating the PubMedBERT language model with a ViT, which is followed by BiomedCLIP generating visual and textual inputs for better classification. They also performed random transformations to generate augmented data for training.
@@
Team WueVision used a domain-adaptive pre-training strategy with the EVA-02 model on ImageNET and their custom dataset, EndoExtend24, a fusion of ten public and private endoscopy datasets. Class imbalance was mitigated using data augmentations.
@@
Team Llama_Mamba implemented FasterViT-3, an optimized ViT model known for computational efficiency. They adapted the model by modifying the final classification layer to align with the Capsule dataset’s categories and used pre-trained weights for initialization. No explicit class imbalance mitigation techniques or unique methodologies were mentioned.
@@
Team Seq2Cure proposed a multi-model ensemble combining CNN and transformer architectures to leverage complementary strengths. They addressed class imbalance with weighted random sampling, focal loss, and data augmentations.
@@
Team Taaldhwaj proposed a CAVE-Net (Classifying Abnormalities in Video Capsule Endoscopy Network), an ensemble image classification model. It combines predictions from three models using soft voting: a CBAM-enhanced ResNet, a DNN based on a pre-trained ResNet-50 autoencoder, and Syn-XRF, which includes SVM, Random Forest, KNN, and XGBoost. Data augmentation techniques like flips, rotations, zooming, cropping, shearing, elastic transformations, Gaussian noise, and image sharpening were used to balance the dataset and create up to 7,500 samples per class.
@@
Team Capsule Commandos concluded with DaViT (Dual Attention Vision Transformers) with two attention blocks—spatial window attention and channel group attention, as their chosen model for the challenge while also implementing custom CNN, ResNet-50, ViT, and Multiscale ViT on the training and validation dataset.
@@
Team eAI implemented EfficientViT-l2, an optimized ViT model known for computational efficiency. The team used pre-trained weights for initialization and mitigated class imbalance with weighted sampling, weighted cross-entropy loss, and augmentations such as MixUp.
@@
Team VCap proposed a multi-backbone feature extraction approach incorporating ResNet50, DeiT, and MobileNetV3-Large. They employed SMOTE for addressing class imbalance in under-represented classes. Their use of diverse backbone architectures and targeted re-sampling techniques highlighted a unique and balanced strategy.
@@
Team DS&Chill implemented a hybrid ensemble model combining YOLOv11 for object detection and DenseNet121 for classification, leveraging a soft voting mechanism to integrate predictions from both models. They addressed class imbalance by augmenting the dataset by merging the training and validation dataset and re-splitting it along with targeted transformations.
@@
Team Layer Players incorporated Squeeze and Excitation blocks into the ResNet-101 architecture for training using cross-entropy loss and Adam optimizer with learning rate scheduler for optimized performance. They have performed standard preprocessing techniques of normalization and random transformations.
Team Rookies implemented a Swin Transformer pre-trained on ImageNet. Their report did not specify any class imbalance mitigation techniques.
@@
Team Team_CSIR explored multiple models, including ResNet-18, ResNet-50, SenseNet-121, DenseNet-169, and MobileNetV3, and concluded MobileNetV3 with a self-attention mechanism as it showed the best performance. The issue of class imbalance was addressed by implementing focal loss.
@@
Team CapsuleNet fine-tuned EfficientNet-b7 with PReLU activation functions for better feature representation. Selective class-based augmentations ensured approximately 1500 samples per class to mitigate the imbalance. The use of PReLU activation introduced a unique optimization in learning non-linear features.
@@
Team Code Cortex developed a CASCRNet (Capsule Endoscopy-ASPP-SCR-Network), a hybrid model combining atrous spatial pyramid pooling blocks with shared channel residual blocks. No specific class imbalance mitigation techniques were mentioned. The use of pooling blocks for efficient feature extraction added depth to their methodology.
@@
Team ViFo Tech implemented a hybrid ResNet-50 and SVM model. They handled class imbalance using data augmentation techniques, including rotations, flips, zooms, and brightness adjustments. Their integration of deep learning with traditional SVM classification offered a distinct perspective.
@@
Team 1b2w implemented transfer learning with Swin Transformer and applied the learning rate scheduler to adjust parameters and prevent overfitting.
@@
Team Machine Minds proposed an Efficient Fusion U-Net that integrated EfficientNet B7 with a U-Net encoder. They addressed class imbalance through Gaussian noise removal.
@@
Team aiVengers employed a Swin Tranformer model combined with self-supervised contrastive learning using a two-phase training strategy of contrastive pre-training followed by supervised fine-tuning, enhancing feature learning from limited labeled data. Class imbalance is mitigated through contrastive learning involving triplet-based training and data augmentation to improve learning for minority classes.
@@
Team Optiminds utilized multiple ViT models, including DeiT, DinoV2, and Swin Transformers, to fine-tune performance. No specific class imbalance techniques were reported.
@@
Team Pioneers utilized a pre-trained ViT model from HuggingFace, fine-tuned on ImageNet-21k. They employed a balanced DataLoader with weighted sampling to address class imbalance. Despite their efforts, their submission did not meet qualification standards.
@@
Team Organic implemented a custom CNN architecture for multi-class classification of VCE frames. This model integrated multiple convolutional layers with varying kernel sizes ($3\times 33 × 3$, $5\times 55 × 5$, $7\times 77 × 7$), max pooling, and strategic dropout application to enhance feature extraction and generalization for anomaly detection. The model was compiled using an SGD optimizer and categorical cross-entropy loss function, achieving low loss scores that indicated no overfitting.
@@
Team Stem Sisters used a semi-supervised approach with transductive learning and GLCM-based texture features for multi-class VCE image classification, outperforming other methods like Random Forest and SVM. They addressed class imbalance through data augmentation for minority classes and under-sampling for overrepresented classes. Their unique use of transductive learning, which focuses on specific un-labelled data, proved effective.
@@
Team DeepScope Innovators employed an ensemble of DenseNet and ResNet for robust feature extraction. The team utilized normalization and augmentation techniques, such as random horizontal flips and rotations, to mitigate class imbalance. Their focus on ensemble methods added robustness to the classification task.
@@
Team BotBotBot implemented a custom ViT-CNN hybrid architecture, combining a ViT branch and a ResNet34 CNN branch, followed by a classification head that performs binary classification first, and then specific abnormality is classified. They applied three levels of augmentations (Heavy, Medium, and Light) based on the intensity of transformations.
@@
Team Deep_Learners implemented FuseCaps, a Fusion-based Framework for Capsule Endoscopy. This hybrid feature extraction method combines CNN, specifically DenseNet, for feature extraction concatenated with MLP, which performed dimensionality reduction on the features extracted using Radiomics on the training dataset. This addressed the class imbalance of the dataset by capturing both deep and handmade representations.
@@
Team EndoAI introduced a ResNet-18-based encoder-decoder framework augmented with omni dimensional gated attention and wavelet transformation modules. These additions aimed to enhance feature localization and extraction. Their integration of wavelet transformations was a unique aspect of their methodology.

### v3

@@ Abstract
We present the Capsule Vision 2024 Challenge: Multi-Class Abnormality Classification for Video Capsule Endoscopy. It was virtually organized by the Research Center for Medical Image Analysis and Artificial Intelligence (MIAAI), Department of Medicine, Danube Private University, Krems, Austria in collaboration with the 9th International Conference on Computer Vision & Image Processing (CVIP 2024) being organized by the Indian Institute of Information Technology, Design and Manufacturing (IIITDM) Kancheepuram, Chennai, India. This document provides an overview of the challenge, including the registration process, rules, submission format, description of the datasets used, qualified team rankings, all team descriptions, and the benchmarking results reported by the organizers.
@@ VI-A Description of all the received entries
Team PuppyOps utilized DINOv2, combining a vision transformer for feature extraction and fully connected layers for multi-class classification. Random flips and brightness adjustments were applied as data augmentation techniques. Link to access code.
@@
Team MedInfoLab IIT Hyderabad performed fine-tuning of BiomedCLIP-PubMedBERT by integrating the PubMedBERT language model with a ViT, which is followed by BiomedCLIP generating visual and textual inputs for better classification. They also performed random transformations to generate augmented data for training. Link to access code.
@@
Team WueVision applied domain-adaptive pre-training on the EVA-02 mode, adapting it for the general task of GIE using their custom dataset, EndoExtend24 - a fusion of ten public and private endoscopy datasets. To address class imbalance, they employed data augmentations and weighted sampling. Link to access code.
@@
Team Llama_Mamba implemented FasterViT-3, an optimized ViT model known for computational efficiency. They adapted the model by modifying the final classification layer to align with the Capsule dataset’s categories and used pre-trained weights for initialization. No explicit class imbalance mitigation techniques or unique methodologies were mentioned. Link to access code.
@@
Team Seq2Cure proposed a multi-model ensemble combining CNN and transformer architectures, leveraging complementary strengths through full parameter fine-tuning for all models. They addressed class imbalance with weighted random sampling, focal loss, and data augmentations. Link to access code.
@@
Team Taaldhwaj proposed a CAVE-Net (Classifying Abnormalities in Video Capsule Endoscopy Network), an ensemble image classification model. It combines predictions from three models using soft voting: a CBAM-enhanced ResNet, a DNN based on a pre-trained ResNet-50 autoencoder, and Syn-XRF, which includes SVM, Random Forest, KNN, and XGBoost. Data augmentation techniques like flips, rotations, zooming, cropping, shearing, elastic transformations, Gaussian noise, and image sharpening were used to balance the dataset and create up to 7,500 samples per class. Link to access code.
@@
Team Capsule Commandos concluded with DaViT (Dual Attention Vision Transformers) with two attention blocks—spatial window attention and channel group attention, as their chosen model for the challenge while also implementing custom CNN, ResNet-50, ViT, and Multiscale ViT on the training and validation dataset. Link to access code.
@@
Team eAI implemented EfficientViT-l2, an optimized ViT model known for computational efficiency. The team used pre-trained weights for initialization and mitigated class imbalance with weighted sampling, weighted cross-entropy loss, and augmentations such as MixUp. Link to access code.
@@
Team VCap proposed a multi-backbone feature extraction approach incorporating ResNet50, DeiT, and MobileNetV3-Large. They employed SMOTE for addressing class imbalance in under-represented classes. Their use of diverse backbone architectures and targeted re-sampling techniques highlighted a unique and balanced strategy. Link to access code.
@@
Team DS&Chill implemented a hybrid ensemble model combining YOLOv11 for object detection and DenseNet121 for classification, leveraging a soft voting mechanism to integrate predictions from both models. They addressed class imbalance by augmenting the dataset by merging the training and validation dataset and re-splitting it along with targeted transformations. Link to access code.
@@
Team Layer Players incorporated Squeeze and Excitation blocks into the ResNet-101 architecture for training using cross-entropy loss and Adam optimizer with learning rate scheduler for optimized performance. They have performed standard preprocessing techniques of normalization and random transformations. Link to access code.
Team Rookies implemented a Swin Transformer pre-trained on ImageNet for classifying gastrointestinal abnormalities in VCE images. While the report did not explicitly mention specific class imbalance mitigation techniques, the use of a pre-trained Swin Transformer inherently helped the model leverage learned features from large, diverse datasets like ImageNet, which can improve generalization and reduce overfitting. Furthermore, the team employed data augmentation techniques, such as random horizontal flips and image normalization, to increase the diversity of the training data, which could have helped mitigate the effects of any class imbalance by enhancing the model’s robustness across different gastrointestinal conditions. These techniques, alongside the pre-trained model, contributed to the model’s overall performance despite the potential challenges posed by class imbalance. Link to access code.
@@
Team_CSIR explored multiple models, including ResNet-18, ResNet-50, DenseNet-121, DenseNet-169, and MobileNetV3, and concluded MobileNetV3 with a self-attention mechanism as it showed the best performance. The issue of class imbalance was addressed by implementing focal loss. Link to access code.
@@
Team CapsuleNet fine-tuned EfficientNet-b7 with PReLU activation functions for better feature representation. Selective class-based augmentations ensured approximately 1500 samples per class to mitigate the imbalance. The use of PReLU activation introduced a unique optimization in learning non-linear features. Link to access code.
@@
Team Code Cortex developed a CASCRNet (Capsule Endoscopy-ASPP-SCR-Network), a hybrid model combining atrous spatial pyramid pooling blocks with shared channel residual blocks. The team used focal loss to mitigate class imbalance. The use of pooling blocks for efficient feature extraction added depth to their methodology. Link to access code.
@@
Team ViFo Tech implemented a hybrid ResNet-50 and SVM model. They handled class imbalance using data augmentation techniques, including rotations, flips, zooms, and brightness adjustments. Their integration of deep learning with traditional SVM classification offered a distinct perspective. Link to access code.
@@
Team 1b2w implemented transfer learning with Swin Transformer and applied the learning rate scheduler to adjust parameters and prevent overfitting. Link to access code.
@@
Team Machine Minds proposed an Efficient Fusion U-Net that integrated EfficientNet B7 with a U-Net encoder. They addressed class imbalance through Gaussian noise removal. Link to access code.
@@
Team aiVengers employed a Swin Tranformer model combined with self-supervised contrastive learning using a two-phase training strategy of contrastive pre-training followed by supervised fine-tuning, enhancing feature learning from limited labeled data. Class imbalance is mitigated through contrastive learning involving triplet-based training and data augmentation to improve learning for minority classes. Link to access code.
@@
Team Optiminds utilized multiple ViT models, including DeiT, DinoV2, and Swin Transformers, to fine-tune performance. No specific class imbalance techniques were reported. Link to access code.
@@
Team Pioneers utilized a pre-trained ViT model from HuggingFace, fine-tuned on ImageNet-21k. They employed a balanced DataLoader with weighted sampling to address class imbalance. Despite their efforts, their submission did not meet qualification standards. Link to access code.
@@
Team Organic implemented a custom CNN architecture for multi-class classification of VCE frames. This model integrated multiple convolutional layers with varying kernel sizes ($3\times 33 × 3$, $5\times 55 × 5$, $7\times 77 × 7$), max pooling, and strategic dropout application to enhance feature extraction and generalization for anomaly detection. The model was compiled using an SGD optimizer and categorical cross-entropy loss function, achieving low loss scores that indicated no overfitting. Link to access code.
@@
Team Stem Sisters used a semi-supervised approach with transductive learning and GLCM-based texture features for multi-class VCE image classification, outperforming other methods like Random Forest and SVM. They addressed class imbalance through data augmentation for minority classes and under-sampling for overrepresented classes. Their unique use of transductive learning, which focuses on specific un-labelled data, proved effective. Link to access code.
@@
Team DeepScope Innovators employed an ensemble of DenseNet and ResNet for robust feature extraction. The team utilized normalization and augmentation techniques, such as random horizontal flips and rotations, to mitigate class imbalance. Their focus on ensemble methods added robustness to the classification task. Link to access code.
@@
Team BotBotBot implemented a custom ViT-CNN hybrid architecture, combining a ViT branch and a ResNet34 CNN branch, followed by a classification head that performs binary classification first, and then specific abnormality is classified. They applied three levels of augmentations (Heavy, Medium, and Light) based on the intensity of transformations. Link to access code.
@@
Team Deep_Learners implemented FuseCaps, a Fusion-based Framework for Capsule Endoscopy. This hybrid feature extraction method combines CNN, specifically DenseNet, for feature extraction concatenated with MLP, which performed dimensionality reduction on the features extracted using Radiomics on the training dataset. This addressed the class imbalance of the dataset by capturing both deep and handmade representations. Link to access code.
@@
Team EndoAI introduced a ResNet-18-based encoder-decoder framework augmented with omni dimensional gated attention and wavelet transformation modules. These additions aimed to enhance feature localization and extraction. Their integration of wavelet transformations was a unique aspect of their methodology. Link to access code.